Find the atomic mass unit of Sulfur-32 with a mass of 3972 and percent abundance 94.99%,...

  1. Find the atomic mass unit of Sulfur-32 with a mass of 3972 and percent abundance 94.99%, Sulfur-33 with a mass of 32.971 and percent abundance 0.75%, Sulfur-34 with a mass of 33.968 and percent abundance 4.25%, and Sulfur-36 with a mass of 35.967 and percent abundance 0.01%.
  2. Find the atomic mass unit of Silicon-28, Silicon-29, Silicon-30 with masses of 27.977, 28.976, 29.974, respectively. The percent abundance of the silicon isotopes is 92.2%, 4.7% and 3.1%, respectively.
  3. Find the mass ratios and atomic ratios of the following compounds.

  4. PCI3
  5. SbF5
  6. BaCl2
  7. BaI2
  8. CCl4
  9. OsO4
  10. Instructions: Write the formula for each of the following ionic compounds.

  11. Sodium chloride
  12. Potassium chloride
  13. Ammonium chloride
  14. Sodium hydroxide

Would be very helpful to understand how these formulas break down
Answer #1

Formula fox given ionic compounds; O Sodium Chloride - Nac! - Nacio Potassium chloride - Kata k® cp@ * Ammonium chloxice : N4c) Bacl₂ = Ba = 137.327 (1 = 35. 453 Bach, = {1x137. 397) + (2x35.453) = 137.327 + 70.906 - 208. 233 for Ba=0.6595 = 65.95% df) Oson : Os = 190.23 0 3 15.999 Csą, = (1x 190.23) + ( 4x 15.999) = 190-23 + 64 = 254.23 for Os = 0.7452 : 74.82% 0 -0.2517